OpenStudy (angelicahicks29):

Prove this (tanx)/(1+cosx) + (sinx)/(1-cosx)= cotx+secxcscx

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

probably if you add you get it

OpenStudy (angelicahicks29):

\[\frac{ tanx }{ 1+cosx }+\frac{ sinx }{ 1-cosx }=cotx+secxcscx\]

OpenStudy (angelicahicks29):

Hi Misty

OpenStudy (misty1212):

wanna try it?

OpenStudy (angelicahicks29):

Yes. I'm so confused

OpenStudy (misty1212):

it is loads easier if you replace \(\cos(x)\) by \(a\) and \(\sin(x)\) by \(b\) since 98% of this is pure algebra

OpenStudy (misty1212):

so the left hand side, the side we can add is \[\frac{ \tan(x) }{ 1+\cos(x) }+\frac{ \sin(x) }{ 1-\cos(x)}\]

OpenStudy (misty1212):

if we use the letters \(a\) and \(b\) it is easier to see what to do \[\frac{\frac{b}{a}}{1+a}+\frac{b}{1-a}\]

OpenStudy (misty1212):

you see what i did there?

OpenStudy (angelicahicks29):

Yes i'm gonna write it down and then send you my answer

OpenStudy (misty1212):

ok go ahead i am hoping you recognize that the denominator will be \((1+a)(1-a)=1-a^2\)

OpenStudy (misty1212):

how you doing?

OpenStudy (angelicahicks29):

I got it but it wasn't the way you told me cause I couldnt get the concept but I got down to the answer with a whole bunch of work

OpenStudy (angelicahicks29):

tan(x)/[1 + cos(x)] + sin(x)/[1 - cos(x)] = tan(x)[1 - cos(x)]/{[1 + cos(x)][1 - cos(x)]} + sin(x)[1 + cos(x)]/{[1 - cos(x)][1 + cos(x)]} = tan(x)[1 - cos(x)]/[1 - cos^2(x)] + sin(x)[1 + cos(x)]/[1 - cos^2(x)] = tan(x)[1 - cos(x)]/sin^2(x) + sin(x)[1 + cos(x)]/sin^2(x) = {tan(x)[1 - cos(x) + sin(x)[1 + cos(x)]}/sin^2(x) = [tan(x) - tan(x)cos(x) + sin(x) + sin(x)cos(x)]/sin^2(x) = [tan(x) - {sin(x)/cos(x)}cos(x) + sin(x) + sin(x)cos(x)]/sin^2(x) = [tan(x) - sin(x) + sin(x) + sin(x)cos(x)]/sin^2(x) = [tan(x) + sin(x)cos(x)]/sin^2(x) = [{sin(x)/cos(x)} + sin(x)cos(x)]/sin^2(x) = [{1/cos(x)} + cos(x)]/sin(x) = [sec(x) + cos(x)]/sin(x) = sec(x)/sin(x) + cos(x)/sin(x) = sec(x)csc(x) + cot(x) = cot(x) + sec(x)csc(x)